# UCM-Net: A Lightweight and Efficient Solution for Skin Lesion Segmentation using MLP and CNN

### v1

@@ Abstract
Skin cancer is a significant public health problem, and computer-aided diagnosis can help to prevent and treat it. A crucial step for computer-aided diagnosis is accurately segmenting skin lesions in images, which allows for lesion detection, classification, and analysis. However, this task is challenging due to the diverse characteristics of lesions, such as appearance, shape, size, color, texture, and location, as well as image quality issues like noise, artifacts, and occlusions. Deep learning models have recently been applied to skin lesion segmentation, but they have high parameter counts and computational demands, making them unsuitable for mobile health applications. To address this challenge, we propose UCM-Net, a novel, efficient, and lightweight solution that integrates Multi-Layer Perceptions (MLP) and Convolutional Neural Networks (CNN). Unlike conventional UNet architectures, our UCMNet-Block reduces parameter overhead and enhances UCM-Net’s learning capabilities, leading to robust segmentation performance. We validate UCM-Net’s competitiveness through extensive experiments on isic2017 and isic2018 datasets. Remarkably, UCM-Net has less than 50KB parameters and less than 0.05 Giga-Operations Per Second (GLOPs), setting a new possible standard for efficiency in skin lesion segmentation. The source code will be publicly available.
@@ 1 Introduction
    Improved Segmentation: UCM-Net’s segmentation performance is evaluated using mean Intersection over Union (mIoU) and mean Dice similarity score (mDice). On the Isic2017 and Isic2018 datasets, UCM-Net enhances the baseline U-Net model by an average of 3.03% in mIoU and 1.72% in mDice. Notably, UCM-Net outperforms the state-of-the-art EGE-UNet on ISIC 2017 and ISIC 2018 datasets, with respective mean IoU scores of 81.43% (UCM-Net) vs 80.95% and 80.71% (UCM-Net) vs 80.00%.
@@ 2 Related works
Supervised Methods of Segmentation As the technique evolves and develops, the solution of AI for medical image segmentation is from purely applying a convolution neural network(CNN) such as U-Net and Att-UNet [31] to a hybrid structure method like TransFuse [32] and SANet [33]. U-Net is a earest CNN solution on biomedical image segmentation, which replaces pooling operators by upsampling operators. Att-UNet is developed on the top of U-Net adding attention structures. TransFuse is a novel approach that combines Transformers and CNNs with late fusion for medical image segmentation, achieving a strong performance while maintaining high efficiency, with potential applications in various medical-related tasks. SANet [34], the Shallow Attention Network, addresses challenges in polyp segmentation by mitigating color inconsistencies, preserving small polyps through shallow attention modules, and balancing pixel distributions, achieving remarkable performance improvements. Swin-UNet is a UNet-like pure Transformer for medical image segmentation that proposed shifted windows as the encoder to extract context features, and a transformer-based decoder with patch expanding layer performs the up-sampling operation to restore the spatial resolution of the feature maps. MedT [35] is also a transformer-based network architecture, a gated axial-attention model that introduces an additional control mechanism in the self-attention module. ConvUNeXt [36], an efficient model inspired by ConvNeXts [37] and based on the classic UNet architecture, achieving excellent medical image segmentation results with a significantly reduced parameter count while incorporating features such as large convolution kernels, depth-wise separable convolution, residual connections, and a lightweight attention mechanism. UNeXt [38] is introduced as an efficient Convolutional multilayer perceptron (MLP) based network that reduces parameters and computational complexity while achieving superior segmentation performance through tokenized MLP blocks and channel shifting, making it suitable for point-of-care applications. MALUNet [39] and its extended version EGE-UNet [40] develop new attention modules to significantly reduce parameters and computational complexity while achieving powerful skin lesion segmentation performance, making it highly suitable for resource-constrained clinical environments.
@@ 3 UCM-Net
Convolution Block In our designing, we contain two different convolution blocks. The difference between conv block 1 and conv block 2 is from the number of kernel size. Small number of kernel size can not only reduce the number of learning weights, but also reduce the computations usage [41]. We set conv block 2’s kernel to 1 $\times$ 1. And we maintain conv block 1’s kernel size to 3 $\times$ 3 normally since fewer the beginning convolution kernel can affect the features entropy performance from input.
@@
UCM-Net Block The pseudocode of UCM-Net Block 1 presents our defined sequence of operations, which is how we combine CNN with MLP(Linear transformation operation) for feature learning. The input feature structure for CNN operation is a four-dimensional structure containing batch, channel, height, and weight. However, the input feature structure for MLP is a three-dimensional structure, which includes batch, channel, and vector. As the pseudocode shows, we implement several feature transformations to service CNN and MLP learning in different layers.
Loss functions In our solution, we selected the group loss function from EGE-UNet [40]. The loss function can calculate the loss from the scaled layer masks in different stages with ground truth masks. Equation 1 and 2 present the stage loss in different stage layer and output loss in the output layer, which calculated by binary cross-entropy (BCE) and dice loss (Dice) components.
$Loss_{Stage}=\text{Bce}(StagePred,Target)+\text{Dice}(StagePred,Target)$ (1)
$Loss_{Output}=\text{Bce}(OutputPred,Target)+\text{Dice}(OutputPred,Target)$ (2)
Equation 3 represents the loss in different stages. Equation 4 presents the group loss that includes different stages loss and output loss. $\lambda_{i}$ is the weight for different stage. In this paper, we set $\lambda_{i}$ to 0.1, 0.2, 0.3, 0.4 and 0.5 based on i-th stage as shown in Figure 1.
$Loss_{\text{Stages}}=\sum_{i=1}^{5}\lambda_{i}\times Loss_{\text{Stagei}}$ (3)
$GroupLoss$ $=Loss_{\text{Output}}+Loss_{\text{Stages}}$ (4)
@@ 4.1 Experiments Setting
Datasets To evaluate the efficiency and performance of model with other published models, we pick the two public skin segementation datasets from International Skin Imaging Collaboration, namely ISIC 2017 and ISIC2018. we select two public skin lesion segmentation datasets, namely ISIC2017 [42, 43] and ISIC2018 [44, 45]. The ISIC2017 dataset comprises 2150 dermoscopy images, and ISIC2018 includes 2694 images. We noted that earlier studies[40, 39] have already presented a dataset version with a pre-established train-test partition, maintaining a 7:3 ratio. In our experimental setup, we opted to utilize the previously published dataset version.
Implementation Details Our UCM-Net is implemented with Pytorch [46] framework. All experiments are conducted on the instance node at Lambda [47] that has a single NVIDIA RTX A6000 GPU (24 GB), 14vCPUs, 46 GiB RAM and 512 GiB SSD. The images are normalized and resized to 256$\times$256. Simple data augmentations are applied, including horizontal flipping, vertical flipping, and random rotation. We noticed the prior studies [40, 39] applied initial image processing with the calculated mean and standard deviation (std) values of the whole train and test datasets separately. While this approach can potentially enhance their models’ training and testing performance, the outcomes are notably influenced by the computed mean and std values. Additionally, if the test dataset’s context information is unknown, this operation can render the trained model less practical. In our experiment, we don’t calculate the mean and std values based on the train and test datasets. Besides, TransFuse-S and SwinNet require the pre-train models with the specified input image size in their encoding stage. To enable fair benchmark testing, we follow the image input size for the TransFuse-S [32, 48] and SwinNet [33, 49] to 192$\times$256 and 224$\times$224, Correspondingly. For the optimizer, we select AdamW [50] initialized with a learning rate of 0.001 and a weight decay of 0.01. The CosineAnnealingLR [51] is Utilized as the scheduler with a maximum number of iterations of 50 and a minimum learning rate of 1e-5. A total of 300 epochs are trained with a training batch size of 8 and a testing batch size of 1.
@@
Dice=2×intersectionsum of pixels in prediction+sum of pixels in ground $\text{Dice}=\frac{2\times\text{intersection}}{\text{sum of pixels in prediction}+\text{sum of pixels in ground truth}}$ (6)
where intersection represents the number of pixels that are common between the predicted output and the ground truth, and union represents the total number of pixels in both the predicted output and the ground truth.
@@ 4.2 Performance Comparisons
Table 1: Comparative prediction results on the ISIC2017 and ISIC2018 dataset
@@
isic2017 U-Net [52, 53] 2015 78.12 ± 0.175 85.97 ± 0.196 76.42 ± 0.381 86.63 ± 0.245
@@
isic2018 U-Net [52, 53] 2015 79.86 ± 0.075 87.57 ± 0.085 78.27 ± 0.300 87.81 ± 0.188
@@
EGE-UNet [40, 60] 2023 80.95 ± 0.054 88.23 ± 0.096 80.11 ± 0.221 88.95 ± 0.137
UCM-Net (ours) 2023 81.43 ± 0.020 88.53 ± 0.024 80.74 ± 0.285 89.34 ± 0.175
  • ${}^{\ast}$: this method needs the pre-train model on training.
@@
Models Year Image Size (H x W) Params↓ GFLOPs↓
U-Net [52, 53] 2015 256 x 256 31,037,633 54.7378
Att-UNet [54, 31] 2018 256 x 256 34,878,573 66.6318
TransFuse-S [32, 48] 2021 192 x 256 26,248,725 8.6462
SANet [34, 55] 2021 256 x 256 23,899,497 5.9983
SwinNet [33, 49] 2021 224 x 224 20,076,204 5.5635
MedT [35, 56] 2021 256 x 256 1,564,202 2.4061
ConvUNeXt [36, 57] 2022 256 x 256 3,505,697 7.2537
UNeXt-S [38, 58] 2022 256 x 256 253,561 0.1038
MALUNet [39, 59] 2022 256 x 256 177,943 0.0830
EGE-UNet [40, 60] 2023 256 x 256 53,374 0.0721
UCM-Net(ours) 2023 256 x 256 49,932 0.0465
@@
Table 1 comprehensively evaluates the performance of our UCM-Net, a novel skin lesion segmentation model, compared to well-established models, using the widely recognized ISIC2017 and ISIC2018 datasets. Introduced in 2023, UCM-Net is a robust and highly competitive solution in this domain. One of the key takeaways from the table is UCM-Net’s ability to outperform EGE-UNet, which had previously held the title of the state-of-the-art model for skin lesion segmentation. Our model achieves superior results across various prediction metrics, emphasizing its advancement in the field and its potential to redefine the standard for accurate skin lesion delineation. Moreover, UCM-Net’s performance is notably competitive even when compared to SwinNet, a model that relies on pre-trained models during training. Table 2 complements this assessment by comparing computational aspects and the number of parameters for various segmentation models. Remarkably, UCM-Net, operating with the same number of channels {8,16,24,32,48,64} and image size, as EGE-UNet, boasts fewer parameters and lower GFLOPs. Additionally, even when compared to TransFuse-S and SwinNet, which operate with smaller image sizes, UCM-Net demonstrates faster computational speed. In Figure 5, we present a visual exhibition of all the models’ segmentation outputs. This figure directly compares our segmentation results, those produced by other methods, and the ground truth, all displayed side by side using representative sample images. Notably, our segmentation results demonstrate a remarkable level of accuracy, closely resembling the ground truth annotations. Tables 1-2 and Figure 5 collectively underscore UCM-Net’s exceptional performance and efficiency in skin lesion segmentation, affirming its potential to make a substantial impact in advancing early skin cancer diagnosis and treatment.

### v4

@@ Abstract
Skin cancer presents a formidable public health challenge, with its incidence expected to rise significantly in the coming decades. Early diagnosis is vital for effective treatment, highlighting the importance of computer-aided diagnosis systems in detecting and managing this disease. A critical task in these systems is accurately segmenting skin lesions from images, which is essential for further analysis, classification, and detection. This task is notably complex due to the diverse characteristics of lesions, such as their appearance, shape, size, color, texture, and location, compounded by image quality issues like noise, artifacts, and occlusions. While recent advancements in deep learning have shown promise in skin lesion segmentation, they often need more computational and extensive parameter requirements, limiting their practicality for mobile health applications. Addressing these challenges, we introduce UCM-Net, a novel, efficient, and lightweight model that synergizes the strengths of Multi-Layer Perceptrons (MLP) and Convolutional Neural Networks (CNN). The innovative UCM-Net Block diverges from conventional UNet architectures, offering significant reductions in parameter count and enhancing the model’s learning efficiency. This results in robust segmentation performance while maintaining a low computational footprint. UCM-Net’s performance is rigorously evaluated using the PH2, ISIC 2017, and ISIC 2018 datasets. It demonstrates its effectiveness with fewer than 50K parameters and requires less than 0.05 Giga-Operations Per Second (GLOPs). Moreover, its minimal memory requirement is just 1.19MB in CPU environment positions. It is a potential benchmark for efficiency in skin lesion segmentation, suitable for deployment in resource-constrained settings. In order to facilitate accessibility and further research in the field,the UCM-Net source code is https://github.com/chunyuyuan/UCM-Net.
@@ 1 Introduction
    Light-weight Segmentation: Compared to published models, our model contains lower number of parameters. Besides,we further reduced the model’s memory usage by engineering methods without third party library. Finally, our UCM-Net only spent 1.19MB on CPU environment.
  4. 4.
    New Group Loss Function: In the training parts, we present a new group loss function with the output loss and internal stage losses. The new designed function can improve UCM-Net’s learning ability.
  5. 5.
    Improved Segmentation: UCM-Net’s segmentation performance is evaluated using mean Intersection over Union (mIoU) and mean Dice similarity score (mDice). On the PH2, Isic2017 and Isic2018 datasets, UCM-Net enhances the baseline U-Net model by an average of 2.53% in mIoU and 1.50% in mDice. Notably, UCM-Net outperforms the state-of-the-art EGE-UNet on Ph2, ISIC 2017 and ISIC 2018 datasets, with respective mean IoU scores of 89.62% (UCM-Net) vs 88.39% , 80.71% (UCM-Net) vs 80.22% and 81.26% (UCM-Net) vs 81.16%.
@@ 2 Related works
TinyML for healthcare Integrating the potential of TinyML into healthcare, particularly for tasks such as lesion segmentation, offers an exciting avenue for future research and practical applications. TinyML, which refers to the deployment of machine learning models on low-power, miniature hardware, presents a unique opportunity to bring advanced analytical capabilities directly to the point of care. This approach could democratize access to sophisticated medical image analysis technologies, enabling real-time, on-device processing in environments where traditional computing resources are scarce or in mobile healthcare applications. For instance, applying TinyML for lesion segmentation could facilitate immediate diagnostic insights during patient examinations or in remote areas, significantly reducing the time and infrastructure typically required for such analyses. The integration of TinyML into healthcare devices promises to enhance diagnostic processes, improve patient outcomes, and extend the reach of advanced medical technologies to underserved regions.To further enhance the efficiency and feasibility of deploying TinyML in such critical applications, researchers are exploring innovative approaches like hyper-structure optimization[31] and the use of quantitative methods such as binary neural networks[32]. Hyper-structure optimization aims to minimize the number of parameters without compromising the model’s performance, thus ensuring that the models are lightweight yet effective enough for deployment on tiny devices. As we explore the optimization and application of UCM-Net in this context, we acknowledge the pioneering work and insights offered by Partha. in their exploration of TinyML’s potential in healthcare [33], which serves as a survey foundational reference for our proposed direction.
@@
Supervised Methods of Segmentation As the technique evolves and develops, the solution of AI for medical image segmentation is from purely applying a convolution neural network(CNN) such as U-Net and Att-UNet [34] to a hybrid structure method like TransUNet [35], TransFuse [36] and SANet [37]. U-Net is a earest CNN solution on biomedical image segmentation, which replaces pooling operators by upsampling operators. Att-UNet is developed on the top of U-Net adding attention structures. TransUNet merits both Transformers and U-Net for medical image segmentation. TransFuse is a novel approach that combines Transformers and CNNs with late fusion for medical image segmentation, achieving a strong performance while maintaining high efficiency, with potential applications in various medical-related tasks. SANet [38], the Shallow Attention Network, addresses challenges in polyp segmentation by mitigating color inconsistencies, preserving small polyps through shallow attention modules, and balancing pixel distributions, achieving remarkable performance improvements. Swin-UNet is a UNet-like pure Transformer for medical image segmentation that proposed shifted windows as the encoder to extract context features, and a transformer-based decoder with patch expanding layer performs the up-sampling operation to restore the spatial resolution of the feature maps. MedT [39] is also a transformer-based network architecture, a gated axial-attention model that introduces an additional control mechanism in the self-attention module. ConvUNeXt [40], an efficient model inspired by ConvNeXts [41] and based on the classic UNet architecture, achieving excellent medical image segmentation results with a significantly reduced parameter count while incorporating features such as large convolution kernels, depth-wise separable convolution, residual connections, and a lightweight attention mechanism. UNeXt [42] is introduced as an efficient Convolutional multilayer perceptron (MLP) based network that reduces parameters and computational complexity while achieving superior segmentation performance through tokenized MLP blocks and channel shifting, making it suitable for point-of-care applications. MALUNet [43] and its extended version EGE-UNet [44] develop new attention modules to significantly reduce parameters and computational complexity while achieving powerful skin lesion segmentation performance, making it highly suitable for resource-constrained clinical environments.
@@ 3 UCM-Net
Convolution Block In our designing, we contain two different convolution blocks. Convolution Block 1 utilizes a kernel size of 3×3, which is commonly employed for its effectiveness in capturing spatial relationships within the input features. This size is particularly advantageous in the network’s initial layers, where preserving the spatial integrity of feature maps is essential for decoding complex input patterns. Convolution Block 2 is characterized by a 1×1 kernel size, drastically reducing the number of learnable parameters and computational load[45]. The smaller kernel size serves to linearly combine features across channels, allowing for dimensionality reduction and feature recombination without spatial aggregation, thus enhancing the network’s efficiency.
@@
UCM-Net Block The UCM-Net Block exemplifies a sophisticated approach to combining Convolutional Neural Networks (CNNs) with Multilayer Perceptrons (MLPs) for robust feature learning. This hybrid model benefits from the spatial feature extraction capabilities of CNNs and the pattern recognition strengths of MLPs through a series of strategically designed operations. Initially, the input feature map undergoes reshaping to cater to the distinct structural requirements of CNNs and MLPs—transitioning from a four-dimensional tensor for CNN processing to a three-dimensional tensor for MLP operations. Following this, the block applies a sequence of linear and convolutional transformations, each accompanied by normalization and activation functions. This includes the strategic use of different kernel sizes in the convolutional stages to balance computational efficiency with the network’s learning capacity. Furthermore, a residual connection is integrated, enhancing information flow through the network and supporting the construction of deeper architectures by mitigating the vanishing gradient problem. Through these mechanisms, UCM-Net adeptly captures and processes complex patterns in image data, demonstrating its potential for advanced image analysis applications. The pseudocode of UCM-Net Block 1 presents our defined sequence of operations, which is how we combine CNN with MLP(Linear transformation operation) for feature learning.
Loss functions: In our solution, we designed a group loss function similar to those used in TransFuse [36] and EGE-UNet [44]. However, the base loss function for each stage is distinct. The base loss function from TransFuse and EGE-UNet is calculated using binary cross-entropy (BCE) (1) and Dice loss (Dice) (2) components. Our proposed base loss function is calculated from BCE components and squared-Dice loss (SDice) components, to calculate the loss from the scaled layer masks in different stages compared with the ground truth masks. Equations (1) and (2) present the stage loss in different layers and the output loss in the output layer, which are calculated using binary cross-entropy (BCE) and Dice loss (Dice) components, respectively.
$\hskip 56.9055pt\text{BCE}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right],$ (1)
where $N$ is the total number of pixels (for image segmentation) or elements (for other tasks), $y_{i}$ is the ground truth value, and $p_{i}$ is the predicted probability for the $i$-th element.
[…]
where smooth is a small constant added to improve numerical stability.
[…]
which represents an enhancement over the standard Dice loss by emphasizing the squared terms of intersections and unions.
[…]
Equations (1) and (3) define the base loss functions (5) for our proposed model, incorporating the Dice loss, and squared-Dice loss components. $\lambda_{i}$ is the weight for different stages. In this paper, we set $\lambda_{i}$ to 0.1, 0.2, 0.3, 0.4, and 0.5 based on the i-th stage, as illustrated in Figure 4. Equations (8) is our proposed group loss function that calculates the loss from the scaled layer masks in different stages with ground truth masks. Equation and present the stage loss in different stage layer and output loss in the output layer.
$\hskip 56.9055ptLoss_{Stage}=\text{BCE}(StagePred,Target)+\text{Squared\_Dice\_Loss}(StagePred,Target),$ (6)
$\hskip 56.9055ptLoss_{Output}=\text{BCE}(OutputPred,Target)+\text{Squared\_Dice\_Loss}(OutputPred,Target),$ (7)
$\hskip 56.9055ptGroup\_Loss=Loss_{Output}+\sum_{i=1}^{5}\lambda_{i}\times Loss_{Stage_{i}}$ (8)
@@ 4.1 Experiments Setting
Datasets To evaluate the efficiency and performance of model with other published models, we pick the three public skin segementation datasets from PH2, International Skin Imaging Collaboration, namely ISIC2017 [46, 47] and ISIC2018 [48, 49]. The PH2 [50] is a dermoscopic image database which 400 images of contains manual segmentation and the clinical diagnosis. The ISIC2017 dataset comprises 2150 dermoscopy images, and ISIC2018 includes 2694 images. We noted that earlier studies[44, 43] have already presented a dataset version with a pre-established train-test partition, maintaining a 7:3 ratio. In our experimental setup, we opted to utilize the previously published dataset version.
Implementation Details Our UCM-Net is implemented with Pytorch [51] framework. All experiments are conducted on the instance node at Lambda [52] that has a single NVIDIA RTX A6000 GPU (24 GB), 14vCPUs, 46 GiB RAM and 512 GiB SSD. The images are normalized and resized to 256$\times$256. Simple data augmentations are applied, including horizontal flipping, vertical flipping, and random rotation. We noticed the prior studies [44, 43] applied initial image processing with the calculated mean and standard deviation (std) values of the whole train and test datasets separately. While this approach can potentially enhance their models’ training and testing performance, the outcomes are notably influenced by the computed mean and std values. Additionally, if the test dataset’s context information is unknown, this operation can render the trained model less practical. In our experiment, we don’t calculate the mean and std values based on the train and test datasets. Besides, TransFuse-S and SwinNet require the pre-train models with the specified input image size in their encoding stage. To enable fair benchmark testing, we follow the image input size for the TransFuse-S [36, 53],TransUNet[35, 54] and SwinNet [37, 55] to 192$\times$256, 224$\times$224 and 224$\times$224, Correspondingly. For the optimizer, we select AdamW [56] initialized with a learning rate of 0.001 and a weight decay of 0.01. The CosineAnnealingLR [57] is Utilized as the scheduler with a maximum number of iterations of 50 and a minimum learning rate of 1e-5. A total of 300 epochs are trained with a training batch size of 8 and a testing batch size of 1.
@@
Dice=2×intersection+smooothsum of pixels in prediction+sum of pixels in ground $\text{Dice}=\frac{2\times\text{intersection}+\text{smoooth}}{\text{sum of pixels in prediction}+\text{sum of pixels in ground truth}+\text{smoooth}}$ (10)
where intersection represents the number of pixels that are common between the predicted output and the ground truth, smooth is a small constant added to improve numerical stability and union represents the total number of pixels in both the predicted output and the ground truth.
@@ 4.2 Performance Comparisons
Table 1: Comparative prediction results on the PH2 dataset
@@
PH2 TransUnet [35, 54] 2021 90.77 ± 0.030 95.09 ± 0.014 91.89 ± 0.148 95.77 ± 0.080
SANet [38, 58] 2021 89.99 ± 0.349 94.62 ± 0.206 91.06 ± 0.481 95.32 ± 0.264
SwinNet [37, 55] 2021 88.92 ± 0.493 94.03 ± 0.289 89.96 ± 0.388 94.71 ± 0.215
TransFuse-S [36, 53] 2021 91.15 ± 0.131 95.30 ± 0.078 92.16 ± 0.113 95.92 ± 0.061
U-Net [59, 60] 2015 89.09 ± 0.224 94.13 ± 0.112 90.08 ± 0.545 94.78 ± 0.302
Att-UNet [61, 34] 2018 89.01 ± 0.345 94.06 ± 0.222 89.82 ± 0.371 94.64 ± 0.206
MedT [39, 62] 2021 87.77 ± 0.099 93.35 ± 0.063 88.53 ± 0.338 93.92 ± 0.190
ConvUNeXt [40, 63] 2022 89.46 ± 0.056 94.33 ± 0.035 90.39 ± 0.172 94.95 ± 0.095
UNeXt-S [42, 64] 2022 89.03 ± 0.246 94.09 ± 0.163 89.689 ± 0.269 94.56 ± 0.149
MALUNet [43, 65] 2022 87.75 ± 0.378 93.35 ± 0.214 88.19 ± 0.329 93.73 ± 0.186
EGE-UNet [44, 66] 2023 88.39 ± 0.303 93.73 ± 0.189 88.93 ± 0.802 94.14 ± 0.450
UCM-Net (ours) 2023 89.62 ± 0.221 94.42 ± 0.124 90.51 ± 0.211 94.96 ± 0.116
  • : this method needs the pre-train model on training.
Table 2: Comparative prediction results on the ISIC2017 dataset
Dataset Models Year mIoU(%)↑ mDice(%)↑ mIoU*(%)↑ mDice*(%)↑
isic2017 TransUnet [35, 54] 2021 81.12 ± 0.152 88.38 ± 0.157 80.02 ± 0.011 88.90 ± 0.113
SANet[38, 58] 2021 79.34 ± 0.023 87.04 ± 0.155 79.34 ± 0.102 88.15 ± 0.060
SwinNet [37, 55] 2021 80.51 ± 0.102 87.81 ± 0.231 80.48 ± 0.134 89.19 ± 0.073
TransFuse-S [36, 53] 2021 80.73 ± 0.312 88.03 ± 0.111 79.87 ± 0.109 88.81 ± 0.058
U-Net [59, 60] 2015 78.12 ± 0.175 85.97 ± 0.196 76.42 ± 0.381 86.63 ± 0.245
@@
  • : this method needs the pre-train model on training.
Table 3: Comparative prediction results on the ISIC2018
Dataset Models Year mIoU(%)↑ mDice(%)↑ mIoU*(%)↑ mDice*(%)↑
isic2018 TransUnet [35, 54] 2021 82.00 ± 0.152 89.11 ± 0.031 81.68 ± 0.511 89.92 ± 0.010
SANet [38, 58] 2021 80.37 ± 0.124 87.87 ± 0.114 79.39 ± 0.135 88.51 ± 0.084
SwinNet [37, 55] 2021 81.41 ± 0.069 88.58 ± 0.019 80.72 ± 0.069 89.33 ± 0.042
TransFuse-S [36, 53] 2021 81.69 ± 0.128 88.82 ± 0.134 81.29 ± 0.084 89.68 ± 0.070
U-Net [59, 60] 2015 79.86 ± 0.075 87.57 ± 0.085 78.27 ± 0.300 87.81 ± 0.188
@@
EGE-UNet [44, 66] 2023 81.16 ± 0.104 88.36 ± 0.086 80.28 ± 0.363 89.06 ± 0.223
UCM-Net (ours) 2023 81.26 ± 0.030 88.48 ± 0.109 80.85 ± 0.251 89.35 ± 0.111
  • : this method needs the pre-train model on training.
@@
Models Year Image Size (H x W) Params↓ GFLOPs↓ Memory (MB)↓
TransUNet [35, 54] 2021 224 x 224 105,277,081 24.7278 402.3614
SANet [38, 58] 2021 256 x 256 23,899,497 5.9983 91.9233
SwinNet [37, 55] 2021 224 x 224 20,076,204 5.5635 77.3503
TransFuse-S [36, 53] 2021 192 x 256 26,248,725 8.6462 100.8809
U-Net [59, 60] 2015 256 x 256 31,037,633 54.7378 119.3992
Att-UNet [61, 34] 2018 256 x 256 34,878,573 66.6318 134.0512
MedT [39, 62] 2021 256 x 256 1,564,202 2.4061 6.9670
ConvUNeXt [40, 63] 2022 256 x 256 3,505,697 7.2537 14.3732
UNeXt-S [42, 64] 2022 256 x 256 253,561 0.1038 1.9673
MALUNet [43, 65] 2022 256 x 256 177,943 0.0830 1.6788
EGE-UNet [44, 66] 2023 256 x 256 53,374 0.0721 1.2036
UCM-Net(ours) 2023 256 x 256 49,932 0.0465 1.1905
  • : this method needs the pre-train model on training.
@@
Table 1-3 comprehensively evaluates the performance of our UCM-Net, a novel skin lesion segmentation model, compared to well-established models, using the widely recognized PH2, ISIC2017 and ISIC2018 datasets. Introduced in 2023, UCM-Net is a robust and highly competitive solution in this domain. One of the key takeaways from the table is UCM-Net’s ability to outperform EGE-UNet, which had previously held the title of the state-of-the-art model for skin lesion segmentation. Our model achieves superior results across various prediction metrics, emphasizing its advancement in the field and its potential to redefine the standard for accurate skin lesion delineation. Moreover, UCM-Net’s performance is notably competitive even when compared to SwinNet, a model that relies on pre-trained models during training. Table 4 complements this assessment by comparing computational aspects and the number of parameters for various segmentation models. Remarkably, UCM-Net, operating with the same number of channels {8,16,24,32,48,64} and image size, as EGE-UNet, boasts fewer parameters and lower GFLOPs. Additionally, even when compared to TransUNet, TransFuse-S and SwinNet, which operate with smaller image sizes, UCM-Net demonstrates faster computational speed. In Figure 5, we present a visual exhibition of all the non-pretrained models’ segmentation outputs. This figure directly compares our segmentation results, those produced by other methods, and the ground truth, all displayed side by side using representative sample images. Notably, our segmentation results demonstrate a remarkable level of accuracy, closely resembling the ground truth annotations. Tables 1-3 and Figure 5 collectively underscore UCM-Net’s exceptional performance and efficiency in skin lesion segmentation, affirming its potential to make a substantial impact in advancing early skin cancer diagnosis and treatment.